NEW PROJECT ANALYSIS

You must evaluate the purchase of a proposed spectrometer for the R&D department. The base price is $230,000, and it would cost another $34,500 to modify the equipment for special use by the firm. The equipment falls into the MACRS 3-year class and would be sold after 3 years for $115,000. The applicable depreciation rates are 33%, 45%, 15%, and 7%. The equipment would require a $12,000 increase in net operating working capital (spare parts inventory). The project would have no effect on revenues, but it should save the firm $31,000 per year in before-tax labor costs. The firm's marginal federal-plus-state tax rate is 40%.

  1. What is the initial investment outlay for the spectrometer, that is, what is the Year 0 project cash flow? Round your answer to the nearest cent. Negative amount should be indicated by a minus sign.
    $
  2. What are the project's annual cash flows in Years 1, 2, and 3? Round your answers to the nearest cent.

    In Year 1 $

    In Year 2 $

    In Year 3 $

  3. If the WACC is 14%, should the spectrometer be purchased?
    -Select-YesNo
Answer #1

a.Initial Investment Outlay = Base Price + Modification cost + Increase in Working Capital

= 230,000+34,500 +12,000

= -$276,500 since outflow

b.Annual Cash Flows:

Year 1

2

3

Savings in Cost

        31,000

        31,000

       31,000

Less: Depreciation

        87,285

      119,025

       39,675

Net Savings

      (56,285)

      (88,025)

        (8,675)

Less: Tax @40%

      (22,514)

      (35,210)

        (3,470)

Income after Tax

      (33,771)

      (52,815)

        (5,205)

Add: Depreciation

        87,285

      119,025

       39,675

Cash Flow

        53,514

        66,210

       34,470

Add: After tax salvage value

       76,406

Recovery of Working capital

       12,000

Cash Flow

        53,514

        66,210

     122,876

Note: Written down value of machine = 264,500*7% = $18,515

Sale Price = $115,000

Gain on Sale = $96,485

Tax on Gain = $38,594

After tax salvage value = 115,000– 38,594= $76,406

c.NPV = Present value of cash inflows – present value of cash outflows

= 53,514*PVF(14%, 1 year) + 66,210*PVF(14%, 2 years) + 122,876*PVF(14%, 3 years) – 276,500

= 53,514*0.877 + 66,210*0.769 + 122,876*0.675 – 276,500

= -$95,711.432

No, should not be purchased (since NPV is negative)